# CHIPS-FF: Evaluating Universal Machine Learning Force Fields for Material Properties

### v1

@@ Abstract
In addition to benchmarking the properties of bulk materials, we utilized the CHIPS-FF workflow for interface calculations. Fig. 3c depicts the parity plot for work of adhesion (Wad) for various material interfaces. The MAE for Wad is computed with respect to “ground truth” data taken from Ref. 109 (from experiment and theory). As seen in Fig. 3c, predictions for Wad are extremely poor for each uMLFF type. This is not surprising due to the fact that none of these uMLFF are trained on interface data. We went on to analyze one of the best performing predictions for an Al(111)Al2O3(001) interface (see red box on Fig. 3c). From here, we performed an in-plane (xy) grid search to find the optimal energy configuration of the interface. In addition, this in-plane scan can be a test of how smooth the potential energy surface is. On the right panel of Fig. 3c we see the results of this interface scan for a number of uMLFFs. We observe a total lack of symmetry in the xy energy scan for ORB, but a smoother potential energy surface for MACE, SevenNet and OMat24 models, which may be attributed to the equivariant architecture of the uMLFF. Although the MAE is extremely high for all uMLFF models for Wad, it is the lowest for ALIGNN-FF. This can be due in part to ALIGNN-FF being trained on vdW-corrected DFT calculations and having a substantial amount of exfoliable materials in the dataset.
@@
We chose to benchmark the performance of uMLFF models for the generation of amorphous Si (a-Si) by performing melt/quench simulations. Fig. 4 depicts a summary of these results, where the radial distribution function (RDF) is shown as a function of pair separation distance. We benchmarked these uMLFF RDF curves against the RDF curve obtained from computationally expensive AIMD simulations for a-Si. The inset of Fig. 4 depicts the MAE of each RDF curve with respect to AIMD. We observe that architectures such as MACE, SevenNet, and OMat24 (eqV2), CHGNet and ALIGNN-FF have excellent accuracy compared to MatGL and ORB. We observe the MAE to be highest for ORB, which is contradictory to the model’s excellent performance for structural optimization and phonons. The final amorphous structures for AIMD, eqV2 (lowest MAE) and ORB (highest MAE) are given in the inset of Fig. 4. The high accuracy of ALIGNN-FF with regards to a-Si is surprising due to the fact that ALIGNN-FF had inferior performance for a number of properties and is an invariant model. A more thorough assessment of how ALIGNN-FF and other uMLFFs can model amorphous materials will be the subject of future work, which will also utilize the CHIPS-FF infrastructure. These results bring into question whether or not equivariance is necessary to achieve accurate results for amorphous materials with uMLFFs, and more rigorous benchmarking will need to be conducted to bring clarity to this.
@@
In addition to testing these uMLFF models on a smaller and more focused dataset, we also tested on a larger and more diverse dataset for force predictions. Unlike total energy, forces are more independent of exchange-correlation functional and the accuracy of force predictions on various datasets can give valuable insight into how a particular model performs. We chose to benchmark each uMLFF on the MLEARN dataset from Ref. 110, which consists of face-centered cubic (Cu, Ni) and body-centered cubic (Li, Mo) metals and diamond group IV semiconductors (Si, Ge) which span a wide range of crystal structures and bonding environments ($\approx$ 200 - 300 data points for each element). These results are depicted in Table 3. In addition, we benchmarked the accuracy of force predictions on very large datasets that were used to train uMLFFs. These datasets included ALIGNN_FF_DB (307,000 used to train ALIGNN-FF), MPF (188,000 used to train M3GNet), and MPTrj (1.58 million used to train CHGNet, MACE, SevenNet and used in the training of ORB and OMat models). These results are depicted in Table 4. Unsurprisingly, superior accuracy is obtained across all datasets when the OMat and ORB models are used to predict forces. The CHIPS-FF package has built in functions to perform these benchmarking calculations on each respective dataset, with a flexible framework to add additional datasets (such as OMat and Alexandria) and new uMLFFs as they develop over time.
@@
Mo Li Si Ni Cu Ge uMLFF Type Error (eV/Å) Time (s) Error (eV/Å) Time (s) Error (eV/Å) Time (s) Error (eV/Å) Time (s) Error (eV/Å) Time (s) Error (eV/Å) Time (s) alignn_ff 0.869 1.2 x 102 0.122 9.5 x 101 0.324 9.5 x 101 0.303 2.1 x 102 0.216 1.9 x 102 0.395 4.5 x 101 chgnet 0.361 9.9 x 101 0.045 1.4 x 102 0.188 3.6 x 101 0.059 3.1 x 102 0.058 2.9 x 102 0.207 3.3 x 101 eqV2_153M_omat 0.096 3.0 x 103 0.004 3.5 x 103 0.012 6.6 x 103 0.009 1.8 x 104 0.006 9.2 x 103 0.031 4.6 x 103 eqV2_31M_omat 0.098 6.9 x 102 0.004 7.9 x 102 0.016 3.2 x 103 0.010 3.4 x 103 0.005 2.1 x 103 0.033 1.1 x 103 eqV2_31M_omat_mp_salex 0.056 6.9 x 102 0.007 7.9 x 102 0.021 2.2 x 103 0.011 2.1 x 103 0.006 3.9 x 103 0.028 1.1 x 103 eqV2_86M_omat 0.094 1.8 x 103 0.004 2.0 x 103 0.014 9.3 x 103 0.009 9.9 x 102 0.006 5.5 x 103 0.032 2.7 x 103 eqV2_86M_omat_mp_salex 0.048 1.8 x 103 0.005 2.0 x 103 0.016 4.9 x 103 0.009 5.5 x 103 0.005 9.7 x 103 0.030 2.7 x 103 mace 0.292 2.6 x 102 0.035 2.8 x 102 0.149 1.9 x 102 0.055 1.8 x 103 0.059 5.5 x 102 0.203 1.9 x 102 mace-alexandria 0.343 5.4 x 102 0.025 6.2 x 102 0.202 6.1 x 102 0.099 1.2 x 103 0.055 1.2 x 103 0.161 6.4 x 102 matgl 0.333 1.0 x 101 0.042 2.1 x 101 0.226 4.1 x 101 0.069 7.4 x 101 0.124 6.5 x 101 0.299 2.2 x 101 matgl-direct 0.300 3.8 x 101 0.031 3.9 x 101 0.214 4.7 x 101 0.114 1.6 x 102 0.054 1.4 x 102 0.260 4.8 x 101 orb-v2 0.130 1.1 x 102 0.018 1.2 x 102 0.068 1.4 x 102 0.026 2.8 x 102 0.015 2.8 x 102 0.060 1.4 x 102 sevennet 0.365 1.9 x 102 0.051 1.4 x 102 0.129 1.4 x 102 0.089 8.4 x 102 0.055 3.7 x 102 0.131 1.4 x 102
@@
ALIGNN_FF_DB (307k) MPF (188k) MPTrj (1.58M) MLFF Type Error (eV/Å) Time (s) Error (eV/Å) Time (s) Error (eV/Å) Time (s) alignn_ff 0.472 2.9 x 104 0.639 4.1 x 104 0.581 3.5 x 105 chgnet 0.088 3.2 x 104 0.101 1.1 x 105 0.060 7.6 x 105 eqV2_153M_omat 0.050 1.5 x 106 - - - - eqV2_31M_omat 0.051 3.0 x 105 0.046 3.5 x 105 - - eqV2_31M_omat_mp_salex 0.055 3.9 x 105 0.032 4.3 x 105 - - eqV2_86M_omat 0.052 9.2 x 105 0.043 9.3 x 105 - - eqV2_86M_omat_mp_salex 0.056 9.2 x 105 - - - - mace 0.104 5.0 x 104 0.125 1.8 x 105 0.059 1.1 x 106 mace-alexandria 0.221 1.3 x 105 0.395 2.7 x 105 - - matgl 0.091 1.1 x 104 0.066 3.4 x 104 0.100 1.4 x 105 matgl-direct 0.104 1.7 x 104 0.095 4.2 x 104 0.118 2.7 x 105 orb-v2 0.062 5.0 x 104 0.049 8.4 x 104 0.028 7.5 x 105 sevennet 0.092 4.4 x 104 0.081 6.0 x 104 0.041 5.0 x 105

### v3

@@ Abstract
In addition to benchmarking the properties of bulk materials, we utilized the CHIPS-FF workflow for interface calculations. Fig. 3c depicts the parity plot for work of adhesion (Wad) for various material interfaces. The MAE for Wad is computed with respect to “ground truth” data taken from Ref. 113 (from experiment and theory). As seen in Fig. 3c, predictions for Wad are extremely poor for each uMLFF type. This is not surprising due to the fact that none of these uMLFF are trained on interface data. We went on to analyze one of the best performing predictions for an Al(111)Al2O3(001) interface (see red box on Fig. 3c). From here, we performed an in-plane (xy) grid search to find the optimal energy configuration of the interface. In addition, this in-plane scan can be a test of how smooth the potential energy surface is. On the right panel of Fig. 3c we see the results of this interface scan for a number of uMLFFs. We observe a smoother potential energy surface for MACE, SevenNet, MatterSim and OMat24 models. Although the MAE is extremely high for all uMLFF models for Wad, it is the lowest for ALIGNN-FF. This can be due in part to ALIGNN-FF being trained on vdW-corrected DFT calculations and having a substantial amount of exfoliable materials in the dataset.
@@
We chose to benchmark the performance of uMLFF models for the generation of amorphous Si (a-Si) by performing melt/quench simulations. Fig. 4 depicts a summary of these results, where the radial distribution function (RDF) is shown as a function of pair separation distance. We benchmarked these uMLFF RDF curves against the RDF curve obtained from computationally expensive AIMD simulations for a-Si. The inset of Fig. 4 depicts the MAE of each RDF curve with respect to AIMD. We observe that architectures such as MACE, SevenNet, and OMat24 (eqV2), CHGNet, ALIGNN-FF, MatterSim and ORB have excellent accuracy, while MatGL has a slightly higher MAE and a lower R2 value. The final amorphous structures for AIMD, MatterSim, eqV2 and ORB (best performing uMLFF models) are given in the inset of Fig. 4. The exceptional performance of ORB and MatterSim (both invariant uMLFFs) for a-Si is significant due to the fact that it is much less costly than other similarly performing equivariant uMLFF models. The relatively high accuracy of ALIGNN-FF with regards to a-Si is surprising due to the fact that ALIGNN-FF had inferior performance for a number of properties and is also an invariant model. A more thorough assessment of how ALIGNN-FF and other uMLFFs can model amorphous materials will be the subject of future work, which will also utilize the CHIPS-FF infrastructure. These results bring into question whether or not equivariance is necessary to achieve accurate results for amorphous materials with uMLFFs, and more rigorous benchmarking will need to be conducted to bring clarity to this.
@@
In addition to testing these uMLFF models on a smaller and more focused dataset, we also tested on a larger and more diverse dataset for force predictions. Unlike total energy, forces are more independent of exchange-correlation functional and the accuracy of force predictions on various datasets can give valuable insight into how a particular model performs. We chose to benchmark each uMLFF on the MLEARN dataset from Ref. 114, which consists of face-centered cubic (Cu, Ni) and body-centered cubic (Li, Mo) metals and diamond group IV semiconductors (Si, Ge) which span a wide range of crystal structures and bonding environments ($\approx$ 200 - 300 data points for each element). These results are depicted in Table 3. In addition, we benchmarked the accuracy of force predictions on very large datasets that were used to train uMLFFs. These datasets included ALIGNN_FF_DB (307,000 used to train ALIGNN-FF), MPF (188,000 used to train M3GNet), and MPTrj (1.58 million used to train CHGNet, MACE, SevenNet and used in the training of ORB and OMat models). These results are depicted in Table 4. Unsurprisingly, superior accuracy is obtained across all datasets when the OMat and ORB models are used to predict forces. We also observe highly accurate forces for MatterSim and a substantial reduction in force error when going from MACE-MP-0 to MACE-MPA-0. The CHIPS-FF package has built in functions to perform these benchmarking calculations on each respective dataset, with a flexible framework to add additional datasets (such as OMat and Alexandria) and new uMLFFs as they develop over time.
@@
Mo Li Si Ni Cu Ge uMLFF Type Error (eV/Å) Time (s) Error (eV/Å) Time (s) Error (eV/Å) Time (s) Error (eV/Å) Time (s) Error (eV/Å) Time (s) Error (eV/Å) Time (s) alignn_ff 0.869 1.2 x 102 0.122 9.5 x 101 0.324 9.5 x 101 0.303 2.1 x 102 0.216 1.9 x 102 0.395 4.5 x 101 chgnet 0.361 9.9 x 101 0.045 1.4 x 102 0.188 3.6 x 101 0.059 3.1 x 102 0.058 2.9 x 102 0.207 3.3 x 101 eqV2_153M_omat 0.096 3.0 x 103 0.004 3.5 x 103 0.012 6.6 x 103 0.009 1.8 x 104 0.006 9.2 x 103 0.031 4.6 x 103 eqV2_31M_omat 0.098 6.9 x 102 0.004 7.9 x 102 0.016 3.2 x 103 0.010 3.4 x 103 0.005 2.1 x 103 0.033 1.1 x 103 eqV2_31M_omat_mp_salex 0.056 6.9 x 102 0.007 7.9 x 102 0.021 2.2 x 103 0.011 2.1 x 103 0.006 3.9 x 103 0.028 1.1 x 103 eqV2_86M_omat 0.094 1.8 x 103 0.004 2.0 x 103 0.014 9.3 x 103 0.009 9.9 x 102 0.006 5.5 x 103 0.032 2.7 x 103 eqV2_86M_omat_mp_salex 0.048 1.8 x 103 0.005 2.0 x 103 0.016 4.9 x 103 0.009 5.5 x 103 0.005 9.7 x 103 0.030 2.7 x 103 mace 0.292 2.6 x 102 0.035 2.8 x 102 0.149 1.9 x 102 0.055 1.8 x 103 0.059 5.5 x 102 0.203 1.9 x 102 mace-alexandria 0.343 4.3 x 102 0.025 4.7 x 102 0.202 1.1 x 103 0.099 2.2 x 103 0.055 2.1 x 103 0.161 9.4 x 102 mace-mpa 0.130 2.1 x 102 0.014 2.5 x 102 0.083 3.5 x 102 0.026 7.6 x 102 0.013 7.4 x 102 0.063 3.5 x 102 matgl 0.333 1.0 x 101 0.042 2.1 x 101 0.226 4.1 x 101 0.069 7.4 x 101 0.124 6.5 x 101 0.299 2.2 x 101 matgl-direct 0.300 3.8 x 101 0.031 3.9 x 101 0.214 4.7 x 101 0.114 1.6 x 102 0.054 1.4 x 102 0.260 4.8 x 101 mattersim 0.121 1.2 x 102 0.024 2.2 x 101 0.067 1.5 x 102 0.021 5.3 x 102 0.012 4.6 x 102 0.068 1.3 x 102 orb-v2 0.130 1.1 x 102 0.018 1.2 x 102 0.068 1.4 x 102 0.026 2.8 x 102 0.015 2.8 x 102 0.060 1.4 x 102 sevennet 0.365 1.9 x 102 0.051 1.4 x 102 0.129 1.4 x 102 0.089 8.4 x 102 0.055 3.7 x 102 0.131 1.4 x 102
@@
ALIGNN_FF_DB (307k) MPF (188k) MPTrj (1.58M) MLFF Type Error (eV/Å) Time (s) Error (eV/Å) Time (s) Error (eV/Å) Time (s) alignn_ff 0.472 2.9 x 104 0.639 4.1 x 104 0.581 3.5 x 105 chgnet 0.088 3.2 x 104 0.101 1.1 x 105 0.060 7.6 x 105 eqV2_153M_omat 0.050 1.5 x 106 - - - - eqV2_31M_omat 0.051 3.0 x 105 0.046 3.5 x 105 - - eqV2_31M_omat_mp_salex 0.055 3.9 x 105 0.032 4.3 x 105 - - eqV2_86M_omat 0.052 9.2 x 105 0.043 9.3 x 105 - - eqV2_86M_omat_mp_salex 0.056 9.2 x 105 - - - - mace 0.104 5.0 x 104 0.125 1.8 x 105 0.059 1.1 x 106 mace-alexandria 0.221 1.3 x 105 0.395 2.7 x 105 - - mace-mpa 0.072 7.0 x 105 0.061 1.3 x 105 - - matgl 0.091 1.1 x 104 0.066 3.4 x 104 0.100 1.4 x 105 matgl-direct 0.104 1.7 x 104 0.095 4.2 x 104 0.118 2.7 x 105 mattersim 0.071 8.7 x 104 0.047 7.1 x 104 0.059 1.7 x 106 orb-v2 0.062 5.0 x 104 0.049 8.4 x 104 0.028 7.5 x 105 sevennet 0.092 4.4 x 104 0.081 6.0 x 104 0.041 5.0 x 105